A uniform wire of resistance 4 ohm is bent into a circle of radius r. A specimen of the same wire is connected along the diameter of the circle. What is the eq resistance?

Grade:12

3 Answers

Umakant biswal
7 years ago
@ surabhi 
i think the data given is insuficient to calculate the equivalent resistance . 
its not mentioned that the speciman wire have resistance or not . 
so, i am assuming it to be present of 4 ohm of resister . 
if so, thenit will be parallel combination of two resisters of4 ohm each /. 
so,, acc to the formula 
1/ R (EQUI ) = 1/R1 + 1/R2 
SO, 1/R = 1/4+1/4
1/R= 2/4 = ½  
so, r = 2 ohm 
HOPE U ARE CLEARS NOW 
IF ANSWER ISGIVEN SOMETHING DIFFRENT THEN LETME KNOW 
ALL THE BEST ..
Raman Mishra
7 years ago
Let the resistance per unit length be \lambda. Then,
 \lambda = 4/2\pi r
=> resistance of the specimen wire connected along the diameter is:
R = \lambda \times 2r = \frac{4}{\pi } \Omega
Now after the specimen wire is connected along the diameter, three resistance come in parallel:
2 \Omega (resistance of one half of the circumference),  2 \Omega (resistance of the other half of the circumference) and R(resistance of the diameter). Solving, we get:
{\color{DarkRed}R_{eq} = \frac{4}{4+\pi } \Omega }
Hope you understood dear Surabhi.
RANJEET KUMAR
7 years ago
dear umakant bro..u do not have basic understanding of english neither you know physics well. please do not mislead students if you do not know the answer.